# DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual-Systems

### v1

@@ Abstract
Task oriented dialog systems often rely on static exploration strategies that do not adapt to dynamic dialog contexts, leading to inefficient exploration and suboptimal performance. We propose DyBBT, a novel dialog policy learning framework that formalizes the exploration challenge through a structured cognitive state space $\mathcal{C}$ capturing dialog progression, user uncertainty, and slot dependency. DyBBT proposes a bandit inspired meta-controller that dynamically switches between a fast intuitive inference (System 1) and a slow deliberative reasoner (System 2) based on real-time cognitive states and visitation counts. Extensive experiments on single- and multi-domain benchmarks show that DyBBT achieves state-of-the-art performance in success rate, efficiency, and generalization, with human evaluations confirming its decisions are well aligned with expert judgment. Code is available at https://github.com/carsonz/DyBBT.
@@ 1 Introduction
“The affordances of the environment are what it offers the animal, what it provides or furnishes, for good or ill."
— James J. Gibson, The Ecological Approach to Visual Perception (1979)
@@
Task oriented dialog system (TODS) assist users in achieving specific goals, like booking flights or reserving restaurants, via multi-turn natural language interactions. Dialog policy typically formulated as a sequential decision making problem addressed with Deep Reinforcement Learning (DRL) Nachum et al. (2017); Silver et al. (2014), is bottlenecked by the exploration-exploitation dilemma: balancing exploitation of known rewards against exploration of unknown actions to discover better strategies. Unlike in standard RL, this dilemma in TODS is fundamentally exacerbated by its intrinsic cognitive structure, a dynamic partially observable context characterized by quantifiable features such as the progress ratio of filled goal slots, the entropy of user intent over possible values, and the conditional dependency between unfilled slots based on domain ontology Peng et al. (2017); Wen et al. (2017). These features directly govern the cost benefit analysis of exploration: early in a dialog, high entropy makes information gathering actions valuable; late in a dialog, high slot dependency makes exploitation critical to avoid constraint violations Qin et al. (2023); Zhao et al. (2024).
Exploration in TODS is fundamentally challenging due to its dynamic, partially observable nature Lee et al. (2023), characterized by three key cognitive properties that unfold in distinct dialog phases. Early dialog stages afford information gathering, as user goals are often ambiguous and multiple slots remain unfilled Kwan et al. (2023); Mid-stages afford clarification and confirmation as slots begin to fill and dependencies emerge Jia et al. (2024); and late stages afford task completion, where actions must adhere to strict slot-value dependencies, for example, a taxi cannot be booked without both “departure" and “destination" Niu et al. (2024). This dynamic “affordance landscape" demands adaptive exploration: static strategies cause inefficiencies, premature exploitation fails tasks, while aimless exploration wastes turns.
@@
Current methods for enhancing exploration in TODS, while powerful, are fundamentally misaligned with this dynamic cognitive reality. As illustrated in Figure 1, traditional DRL methods rely on static heuristics like $\epsilon$-greedy Niu et al. (2024), which cannot adapt to the shifting exploration needs across dialog phases. Evolutionary methods like EIERL Zhao et al. (2025) enable global search via population based optimization and elite injection to accelerate evolution, yet struggle in complex multi-domain scenarios due to poor scalability and unflexible updates. LLM based policies Zhang et al. (2024); He et al. (2022) or reasoning techniques such as Tree of Thoughts (ToT) Yao et al. (2023) support deep deliberative planning, but incur prohibitive computational overhead and lack a principled mechanism to trigger such costly reasoning only when necessary. This misalignment reveals a key Research Question: How to design a dialog policy that dynamically perceives cognitive affordances to balance exploration and exploitation?
Figure 1: Traditional DRL methods (left) employ a static exploration strategy with a single policy. EIERL (middle) uses population based evolutionary optimization with elite injection but struggles to scale to complex multi domain tasks. DyBBT (right) introduces a cognitive meta-controller dynamically balances fast intuitive responses and slow deliberative reasoning for adaptive policy selection.
@@
To solve above challenges, we propose DyBBT, a novel framework that grounds decisions in an interpretable cognitive state space $\mathcal{C}$ capturing dialog progress $d_{t}$, user uncertainty $u_{t}$, and slot dependency $p_{t}$, as shown in Figure 1. DyBBT introduces a lightweight meta-controller that dynamically switches between a fast System 1 (for routine decisions) and a slow System 2 (for costly deliberation) based on real-time cognitive signals and visitation counts. This design ensures that expensive reasoning is invoked only when the cognitive state signals under exploration or high uncertainty, addressing the core limitations (RQ) of previous methods. By formalizing dialog affordances and embedding them into a bandit inspired switching mechanism, DyBBT achieves a principled and efficient balance between exploration and exploitation.
@@
In summary, our work makes the following contributions: (1) Formalization of TODS exploration challenge via a structured cognitive state space $\mathcal{C}$ (Section 3.1). (2) Proposal of DyBBT, a novel framework with bandit inspired meta-controller to dynamically balance between fast System 1 and deliberate System 2 reasoning (Section 3.2). (3) Demonstration of state-of-the-art (SOTA) performance and human aligned decisions through extensive experiments (Section 4).
@@ 2 Related Work
2.1 Dialog Policy Learning with Deep Reinforcement Learning
@@
Deep Reinforcement Learning (DRL) has become a dominant paradigm for dialog policy optimization due to its capacity for sequential decision making. Early work applied value based methods Peng et al. (2018) and Policy Gradient Silver et al. (2014) to TODS, Proximal Policy Optimization (PPO) Schulman et al. (2017) was later adopted for improved stability and has become a common baseline. A key limitation of these methods is their reliance on static exploration strategies, such as $\epsilon$-greedy or entropy bonus. These heuristics cannot adapt to the dynamic uncertainty and structural complexity of multi-domain dialogs Kwan et al. (2023); Jia et al. (2024). Recent efforts have incorporated Bayesian reasoning Lee et al. (2023), meta-learning Li et al. (2024); Liang et al. (2025), Cascading RL Du et al. (2024), CB-RL Thoma et al. (2025) to allow more adaptive exploration. While promising, they often lack an explicit and interpretable representation of the internal cognitive dialog state that directly governs exploration, a gap our cognitive state space $\mathcal{C}$ aims to fill.
2.2 Evolutionary and Population Based Methods for Exploration
Evolutionary Reinforcement Learning (ERL) combines population based global search with gradient-based optimization to enhance exploration diversity. Methods such as EIERL Zhao et al. (2025) inject elite policies to accelerate evolution, enabling escape from local optima. However, ERL scales poorly with dialog complexity due to exponential growth in population size Sigaud (2023). Moreover, these methods often rely on fixed schedules for policy replacement, lacking dynamic adaptation to real-time dialog progression and cognitive state changes Bai et al. (2023). In contrast, DyBBT replaces expensive population evolution with a single, efficient dual-system architecture guided by a structured cognitive state space, enabling fine grained, context aware exploration without the scalability limitations of population based approaches.
2.3 Classical and Modern Exploration Theories
The exploration-exploitation trade-off is a cornerstone of sequential decision theory. Bandit algorithms, such as Upper Confidence Bound (UCB) Garivier & Moulines (2011), provide strong theoretical guarantees for stationary settings, and their principles have been extended to contextual bandits Foster & Rakhlin (2020) and hierarchical RL Rohmatillah & Chien (2023). However, directly applying these theories to dialog Partially Observable Markov Decision Processes (POMDPs) faces significant challenges due to non-stationarity, partial observability, and the high dimensional nature of the state space. Our work draws inspiration from the optimism principle of UCB but makes a pragmatic heuristic adaptation to a learned cognitive state space $\mathcal{C}$. This approach preserves the interpretability and theoretical intuition of bandit algorithms while specifically addressing the complexities of sequential dialog environments. Compared to methods like PSRL Chen et al. (2020) that require maintaining a posterior over the entire MDP, our method focuses exploration on a compact cognitive space, offering a computationally efficient alternative better suited to dialog POMDPs.
2.4 Dual-System Architectures and LLMs for dialog Policies
@@
Krämer (2014), combine fast, intuitive processing (System 1) with slow, deliberative reasoning (System 2), have been applied to mathematical reasoning Shi et al. (2024) and commonsense inference Yu et al. (2025). In dialog systems, large language models (LLMs) serve as powerful function approximators Yi et al. (2024), acting as both intuitive generators Ying et al. (2024) and deliberative reasoners Ma et al. (2025). Recent work, such as the Dynamic Dual-Process Transformer He et al. (2024), explicitly models the interaction for dialog policy learning. However, existing switching mechanisms often rely on static heuristics, such as fixed turn counts Qin et al. (2023) or pre-defined confidence thresholds Yao et al. (2023), which lack adaptability and theoretical grounding in exploration. DyBBT addresses this by introducing a meta-controller guided by a bandit inspired principle, dynamically triggering System 2 based on cognitive state visitation counts and parametric uncertainty, offering a principled and efficient alternative to heuristic switching.
@@ 3 Methodology
To address the limitations of static exploration, high computational cost, and lack of cognitive grounding, we propose DyBBT: a framework that grounds adaptive exploration in a structured cognitive state space $\mathcal{C}$, orchestrated by a lightweight meta-controller over a dual-system architecture. Its core innovation is using $\mathcal{C}$ as the decision basis: the meta-controller dynamically triggers costly System 2 based on real-time cognitive signals, enabling efficient, context aware exploration.
@@ 3.1 Theoretical Foundation
3.1.1 Cognitive State Space: A Bridge for Exploration
We define a low dimensional, interpretable representation of the dialog context cognitive state $\mathbf{c}_{t}=[d_{t},u_{t},\rho_{t}]$, to bridge bandit inspired exploration principles with the sequential POMDPs setting. Grounded in Gibson’s affordance theory, $\mathbf{c}_{t}$ captures action possibilities that directly govern the exploration-exploitation trade-off, where $d_{t}$, $u_{t}$, and $\rho_{t}$ quantify dialog progress, user uncertainty, and slot dependency, respectively (see Appendix A.1 for computation).
3.1.2 Reward Smoothness: A Pragmatic Assumption for Structured Tasks
To enable theoretical analysis, we assume the reward function is Lipschitz continuous (Asadi et al., 2018) in the cognitive state space $\mathcal{C}$. We provide empirical validation in Section 5.4.
Assumption 1 (Lipschitz Smooth Reward in $\mathcal{C}$).
@@ 3.1.2 Reward Smoothness: A Pragmatic Assumption for Structured Tasks
3.1.3 Dynamic Balance Principle: A Bandit inspired Heuristic
To address the intractability of optimal exploration in dialog POMDPs, we propose a bandit (Komiyama et al., 2024) inspired heuristic: the exploration bonus for cognitive state $\mathbf{c}_{t}$ is inversely proportional to the square root of its visitation count:
$\text{Exploration-Bonus}(t)\propto\sqrt{\frac{\log T}{n_{t}(\mathbf{c}_{t})}}.$ (1)
This yields sublinear regret: $\mathbb{E}[R(T)]\lesssim\widetilde{\mathcal{O}}\left(L_{r}\cdot\sqrt{\dim(\mathcal{C})\cdot T}\right)$, where $L_{r}$ is the Lipschitz constant. See Appendix A.2 for derivation and assumptions.
Figure 2: The DyBBT Architecture. A meta-controller uses the cognitive state $\mathbf{c}_{t}$, visitation count $n_{t}(\mathbf{c}_{t})$, and System 1’s confidence $p_{t}^{S1}$ to dynamically select between System 1 (fast, intuitive) and System 2 (slow, deliberative). Outputs drive action execution and update visitation/distillation buffers for continuous learning.
@@ 3.2 System Architecture
The proposed DyBBT framework operationalizes its theoretical principles into a dynamic dual-system architecture (Fig. 2), where the meta-controller leverages the cognitive state $\mathbf{c}_{t}$ to orchestrate a cost aware trade-off: invoking the fast, learned System 1 for routine decisions, or the powerful System 2 for high uncertainty or under explored states. This design ensures that expensive deliberation is reserved for contexts where the bandit inspired criterion (Eq. 2) signals its necessity, either due to low visitation ($n_{t}(\mathbf{c}_{t})$) or low confidence ($p_{t}^{S1}$). The result is a closed loop system that translates cognitive perception into adaptive action selection, without redundant or static heuristics.
3.2.1 System 1 (S1): The Fast Intuitive Inference
To provide a low latency, high throughput baseline policy for the majority of dialog turns, mitigating the prohibitive cost of always using a deliberative reasoner, S1 embodies the fast and intuitive system. It is implemented as a base pre-trained Transformer model augmented with LoRA modules Hu et al. (2022). It is first trained via Supervised Fine-Tuning (SFT) on expert trajectories and subsequently optimized using PPO to maximize the expected task reward. At inference time, it takes the current belief state $\mathbf{s}_{t}$, formatted via a concise prompt, and generates an action $a_{t}^{S1}$ along with a calibrated confidence score $p_{t}^{S1}\in[0,1]$, which reflects its intrinsic uncertainty.
3.2.2 System 2 (S2): The Slow Deliberative Reasoner
To handle novel or complex situations where the fast policy (S1) is likely to fail, thereby addressing the suboptimal performance of static DRL policies in under explored regions, S2 represents the slow and analytical system. It utilizes the same base model as S1 but remains frozen to preserve its broad knowledge and reasoning capabilities. The prompt instructs S2 to generate Top-3 distinct action sequences. Each sequence’s quality is evaluated using the ratio of filled key slots. We extract the first action from the highest quality sequence as the output $a_{t}^{S2}$. This system is computationally expensive but is designed to handle novel or high stakes situations identified by the meta-controller.
3.2.3 Meta-Controller: Dynamic Orchestration via Cognitive Affordances
The meta-controller bridges the gap between adaptive exploration and computational efficiency by dynamically invoking S2 only when the cognitive state $\mathbf{c}_{t}$ signals a clear affordance for deliberation. It operates on $\mathbf{c}_{t}=[d_{t},u_{t},\rho_{t}]$, its derived visitation count $n_{t}(\mathbf{c}_{t})$, and S1’s confidence $p_{t}^{S1}$, ensuring decisions are grounded in real time cognitive perception. This hybrid rule combines two practical triggers: (1) exploration in under visited cognitive states, motivated by bandit principles, and (2) uncertainty mitigation when System 1 lacks confidence. The disjunctive design ensures System 2 activation for either systematic exploration or robust decision making.
$\text{Activate S2 IF: }\left(n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}\right)\lor\left(p_{t}^{S1}<\kappa\right).$ (2)
Condition 1: $n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}$, targets under explored regions of $\mathcal{C}$, triggering S2 to perform theoretically motivated exploration where reward potential is high (Assumption 1). This replaces static global heuristics with a local adaptive mechanism. Condition 2: $p_{t}^{S1}<\kappa$, acts as a robustness safeguard, activating S2 when S1 exhibits high parametric or aleatoric uncertainty, often due to partial observability or distributional shift. This prevents catastrophic failures in high stakes states, addressing the brittleness of single policy or fixed schedule methods.
This disjunctive design ensures S2 is invoked only when its cost is justified: to explore promising new territory (curiosity driven) or to mitigate imminent failure (uncertainty driven), achieving both efficiency and robustness. The selected actions are executed via a lightweight action generator ensuring semantic validity. To reduce long term S2 reliance, high quality S2 decisions are distilled into S1 (see Appendix B.5.2 & Algorithm 1).

### v2

# DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual Systems

@@ Abstract
Task oriented dialog systems often rely on static exploration strategies that do not adapt to dynamic dialog contexts, leading to inefficient exploration and suboptimal performance. We propose DyBBT, a novel dialog policy learning framework that formalizes the exploration challenge through a structured cognitive state space $\mathcal{C}$ that captures dialog progression, user uncertainty, and slot dependency. DyBBT proposes a bandit inspired meta-controller that dynamically switches between a fast intuitive inference (System 1) and a slow deliberative reasoner (System 2) based on real-time cognitive states and visitation counts. Extensive experiments on single- and multi-domain benchmarks show that DyBBT achieves SOTA performance in success rate, efficiency, and generalization, with human evaluations confirming that its decisions are well aligned with expert judgment. The code is available at Anonymous Github.
DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual Systems
Shuyu Zhang1thanks: These authors contributed equally to this work.,  Yifan Wei2 *,  Jialuo Yuan3,  Xinru Wang4,  Yanmin Zhu1thanks: Corresponding Author.,   Bin Li5 †  Yujie Liu6, 1Department of Computer Science and Engineering, Shanghai Jiao Tong University 2School of Computer Science and Engineering, Beihang University 3Stanford University 4Data Analytics for Business, University of Sydney 5Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 6School of Information Engineering, Beijing Institute of Graphic Communication carsonz@sjtu.edu.cn, weiyifan@buaa.edu.cn, xinruwang1999@outlook.com yzhu@cs.sjtu.edu.cn, yjl705@stanford.edu, b.li2@siat.ac.cn 18764008899@163.com
@@ 1 Introduction
“The affordances of the environment are what it offers the animal, what it provides or furnishes, for good or ill.”
— The Ecological Approach to Visual Perception (Gibson, 1979)
Figure 1: A comparison of exploration strategies for dialog policy learning. Traditional DRL relies on static heuristics incapable of adapting to dynamic dialog contexts. EIERL uses population based evolution but struggles in complex tasks. DyBBT solves the adaptive exploration challenge by cognitive meta-controller to achieve a principled balance between efficiency and robustness.
@@
Task-oriented dialog systems (TODS) assist users in achieving specific goals through multi-turn interactions. Dialog policy learning is typically formulated as a sequential decision making problem addressed with Deep Reinforcement Learning (DRL) (Nachum et al., 2017; Silver et al., 2014). However, it is fundamentally bottlenecked by the exploration exploitation dilemma: balancing the exploitation of known rewards against the exploration of unknown actions to discover better strategies. In TODS, this dilemma is exacerbated by the dynamic and partially observable context, characterized by quantifiable cognitive features such as dialog progress, user intent entropy, and slot dependencies (Peng et al., 2017; Wen et al., 2017). The features govern the cost benefit of exploration: early in a dialog, high entropy makes information gathering valuable, and later high slot dependency makes exploitation critical to avoid constraint violations (Qin et al., 2023; Zhao et al., 2024).
@@
Existing methods for enhancing exploration in TODS remain misaligned with this dynamic cognitive reality. As illustrated in Fig. 1, traditional DRL methods rely on static heuristics like $\epsilon$-greedy (Niu et al., 2024), which cannot adapt to shifting exploration needs across dialog phases. Evolutionary methods like EIERL (Zhao et al., 2025) enable global search via population based optimization but struggle to scale in complex multi-domain scenarios. LLM-based policies (Zhang et al., 2024; He et al., 2022) or reasoning techniques like Tree of Thoughts (Yao et al., 2023) support deep deliberative planning but incur prohibitive computational overhead and lack a principled mechanism to trigger such costly reasoning only when necessary. This misalignment reveals a key research question: How can we design a dialog policy that dynamically perceives cognitive affordances to balance exploration and exploitation?
@@
Inspired by Gibson’s affordances theory, we propose that the dialog environment presents a dynamic landscape of exploration opportunities which an effective policy must perceive and act upon. We introduce DyBBT, a novel framework that grounds decisions in an interpretable cognitive state space $\mathcal{C}$ that captures dialog progress, user uncertainty, and slot dependency. It employs a lightweight meta-controller that dynamically switches between a fast System 1 for routine decisions and a slow System 2 for costly deliberation, based on real-time cognitive signals. This design ensures expensive reasoning is invoked only when the cognitive state signals high epistemic uncertainty or low confidence, addressing the core limitations of previous methods.
@@
In summary, our contributions are: (1) Formalizing the TODS exploration challenge via a structured cognitive state space $\mathcal{C}$ (Sec. 3.1); (2) Proposing DyBBT, a novel framework with a meta-controller to dynamically balance between fast and slow reasoning (Sec. 3.1.2); (3) Demonstrating SOTA performance and human aligned decisions through extensive experiments (Sec. 4).
@@ 2 Related Work
2.1 DRL for Dialog Policy Learning
@@
Deep Reinforcement Learning (DRL) has become a dominant paradigm for dialog policy optimization due to its capacity for sequential decision making. Early work applied value based methods (Peng et al., 2018) and policy gradient algorithms (Silver et al., 2014) to TODS, with Proximal Policy Optimization (PPO) (Schulman et al., 2017) later adopted for improved stability. A key limitation of these methods is their reliance on static exploration strategies, such as $\epsilon$-greedy or entropy bonus, which cannot adapt to the dynamic uncertainty and structural complexity of multi-domain dialogs (Kwan et al., 2023; Jia et al., 2024). Recent efforts have incorporated Bayesian reasoning (Lee et al., 2023), meta-learning (Li et al., 2024; Liang et al., 2024), and cascading RL (Du et al., 2024) to enable more adaptive exploration. While promising, these approaches often lack an explicit and interpretable representation of the internal dialog state that directly governs exploration, the gap that our structured cognitive state space $\mathcal{C}$ aims to fill.
2.2 Evolutionary Exploration Methods
To overcome static exploration, research has diverged into population based optimization and principled exploration theory. Population based methods like Evolutionary RL (EIERL)  (Zhao et al., 2025) enhance diversity but scale poorly with dialog complexity  (Sigaud, 2023) and lack real time adaptation  (Lin et al., 2025). Theoretically, bandit algorithms (UCB  (Garivier and Moulines, 2011), contextual  (Foster and Rakhlin, 2020), hierarchical RL  (Rohmatillah and Chien, 2023)) and posterior sampling (PSRL  (Chen et al., 2020)) formalize exploration. However, their direct application to dialog POMDPs is hindered by non-stationarity and high dimensionality. Both lines of work lack a mechanism to perceive and respond to the dynamic cognitive affordances such as shifting uncertainty within a dialog. DyBBT bridges this gap by pragmatically adapting bandit inspiration to a learned, low dimensional cognitive state space $\mathcal{C}$, offering a tractable bridge between classical theory and sequential dialog complexity.
2.3 Dual System Reasoning Architectures
@@
Inspired by dual process theory (Krämer, 2014), recent work combines fast processing (System 1) with slow reasoning (System 2) for mathematical reasoning (Shi et al., 2024) and common sense inference (Yu et al., 2025). In TODS, large language models (LLMs) serve as powerful function approximators (Yi et al., 2025), acting as intuitive generators (Ying et al., 2024) and deliberative reasoners (Ma et al., 2025). Frameworks like the Dynamic Dual Process Transformer (He et al., 2024) explicitly model this interaction for policy learning. However, existing switching mechanisms often rely on static heuristics, such as fixed turn counts (Qin et al., 2023) or pre-defined confidence thresholds (Yao et al., 2023), which lack adaptability and theoretical grounding in exploration. DyBBT addresses this by introducing a meta-controller over a cognitive state space, dynamically triggering System 2 based on visitation counts and parametric uncertainty, offering a principled and efficient alternative to heuristic switching.
@@ 3 Methodology
DyBBT formulates dialog exploration as a tractable Contextual Multi-Armed Bandit (CMAB) problem over a cognitive state space $\mathcal{C}$, with theoretical grounding in Lipschitz smooth rewards and sublinear regret. This foundation enables a meta-controller that dynamically triggers System 2 based on visitation counts and confidence scores, balancing exploration and uncertainty in real time.
@@ 3.1 Theoretical Foundation
To provide a principled understanding of how exploration can be efficiently managed in dialog POMDPs, we develop a theoretical analysis that frames the problem as CMAB over a structured cognitive state space $\mathcal{C}$. This formulation rests on three pragmatic foundations: (1) compression of the high dimensional belief state into a low dimensional cognitive representation; (2) a Lipschitz smoothness assumption on the reward function; and (3) a derived bandit style exploration criterion. Together, these steps yield a tractable foundation that directly informs the design of our meta-controller.
3.1.1 CMAB Formulation
We formulate dialog policy learning as a CMAB problem (Foster and Rakhlin, 2020) to render the exploration-exploitation trade-off analytically tractable. The core innovation is a structured cognitive state space $\mathcal{C}$ that compresses the high dimensional belief state into a low dimensional and interpretable representation, thereby bridging bandit theory with dialog POMDPs.
In this CMAB, the arms are $\mathcal{A}={\text{S1},\text{S2}}$ (fast inference System 1 vs. deliberative reasoning System 2). The context is the cognitive state $\mathbf{c}_{t}=[d_{t},u_{t},\rho_{t}]\in\mathcal{C}$ (Computation details in Appendix A.1), which captures dialog progress, user uncertainty, and slot dependency. The reward $r_{t}(a)$ measures task progress and efficiency when choosing arm $a$ in context $\mathbf{c}t$. The objective is to minimize cumulative regret, where $a^{*}_{t}$ is the optimal arm and $a_{t}$ is the chosen arm at turn $t$.
$R_{T}=\sum_{t=1}^{T}\left[\mathbb{E}[r_{t}(a^{*}_{t}\mid\mathbf{c}_{t})]-\mathbb{E}[r_{t}(a_{t}\mid\mathbf{c}_{t})]\right].$ (1)
This formulation treats S2 as an oracle arm. When it invoked, aggressively pursues the optimal action in under explored regions. This directly informs the design of our meta-controller (Sec. 3.2.3).
3.1.2 Reward Smoothness Assumption
To support principled exploration in the cognitive state space $\mathcal{C}$, we assume the reward function is Lipschitz continuous (Asadi et al., 2018; Pazis and Parr, 2013; Ortner and Ryabko, 2012). This standard regularity condition ensures that nearby cognitive states yield similar rewards.
Assumption 3.1 (Lipschitz Smooth Reward in $\mathcal{C}$).
@@ 3.1.2 Reward Smoothness Assumption
This assumption makes visitation counts meaningful proxies for epistemic uncertainty and allows bandit style exploration guarantees to carry over to the dialog POMDP. We empirically validate its practical relevance in Sec. 5.3.
Figure 2: The DyBBT Architecture. A meta-controller uses the cognitive state $\mathbf{c}_{t}$, visitation count $n_{t}(\mathbf{c}_{t})$, and System 1’s confidence $p_{t}^{S1}$ to dynamically select between System 1 (fast intuitive) and System 2 (slow deliberative). Outputs drive action execution and update visitation/distillation buffers for continuous learning.
3.1.3 Dynamic Balance Principle
Building on Assumption 3.1, we derive a principled exploration criterion for the meta-controller by formalizing the exploration-exploitation trade-off as a contextual bandit problem (Kleinberg et al., 2008; Bubeck et al., 2011) The exploration bonus for cognitive state $\mathbf{c}_{t}$ is:
$\text{Exploration-Bonus}(t)\propto\sqrt{\frac{\log T}{n_{t}(\mathbf{c}_{t})}},$ (2)
where $T$ is the total training steps and $n_{t}(\mathbf{c}_{t})$ is the visitation count. This adapts the Upper Confidence Bound (UCB) principle (Ortner and Ryabko, 2012; Foster and Rakhlin, 2020) to the structured cognitive space, with the square root and logarithmic terms arising from concentration inequalities and horizon scaling (Komiyama et al., 2024).
Eq. 2 motivates our meta-controller design. S2 is triggered when the exploration bonus exceeds a threshold, leading to Condition 1: $n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}$. The scalar $\tau$ adjusts the trade-off, analogous to the confidence radius in UCB algorithms.
Under Assumption 3.1 and the approximate MDP structure in $\mathcal{C}$, this exploration strategy achieves sublinear regret (Proof sketch in Appendix A.2), demonstrating efficient and principled exploration in the compressed cognitive space.
@@ 3.2 System Architecture
Building on the theoretical foundation, DyBBT as shown in Fig. 2, operationalizes the CMAB formulation over the cognitive state space $\mathcal{C}$ into a dual-system architecture. The meta-controller directly instantiates the bandit-inspired switching rule (Eq. 2) to dynamically balance between fast intuitive S1 and slow deliberative S2. This principled design ensures expensive S2 is invoked only when cognitive signals and visitation counts indicate high epistemic uncertainty or low confidence, achieving adaptive exploration-exploitation trade-off while maintaining computational efficiency.
3.2.1 S1: Fast Intuitive Inference
S1 serves as the low latency, high throughput policy for most dialog turns, avoiding the prohibitive cost of perpetual deliberation. Given the current belief state, S1 outputs both a system action $a_{t}^{S1}$ formalized as a tuple $\{actiontype,domain,slot\}$, and a confidence score $p_{t}^{S1}\in[0,1]$. This score captures aleatoric uncertainty, complementing the epistemic uncertainty tracked by the meta-controller. S1 is trained in two stages: supervised fine-tuning on expert trajectories to predict actions and calibrated confidence, followed by PPO optimization for task success and efficiency.
3.2.2 S2: Slow Deliberative Reasoner
S2 is invoked only for novel or complex states where S1 is likely to fail. It uses the same frozen base model as S1 to retain broad knowledge. Upon activation, S2 generates three distinct action sequences, evaluates each based on the ratio of filled key slots, and selects the first action of the highest rated sequence as $a_{t}^{S2}$. Although computationally expensive, S2 provides robust reasoning in high uncertainty or high stakes scenarios, as guided by the meta-controller.
3.2.3 Meta-Controller
The meta-controller implements the bandit inspired exploration criterion (Eq. 2) by dynamically selecting between S1 and S2 based on real time cognitive signals. This dual-trigger mechanism bridges bandit theory with practical dialog POMDPs:
Activate S2 If:
$\underbrace{n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}}_{\text{Condition 1}}\;\lor\;\underbrace{p_{t}^{S1}<\kappa}_{\text{Condition 2}}.$ (3)
Condition 1: Exploration. This condition directly implements the exploration bonus from Eq. 2. Under Assumption 3.1, low visitation counts in cognitive region $\mathbf{c}_{t}$ indicate high epistemic uncertainty, justifying systematic exploration via S2. The threshold $\tau\sqrt{\log T}$ adapts the classical bandit confidence radius to our structured cognitive space, ensuring exploration occurs when potential information gain outweighs computational cost.
Condition 2: Confidence. This condition addresses aleatoric uncertainty from partial observability and model limitations. Empirical studies show LLM confidence scores correlate with calibration (Kadavath et al., 2022; Lin et al., 2022; Yin et al., 2023); thus $p_{t}^{S1}<\kappa$ triggers S2 when S1’s parametric knowledge is likely insufficient.
This hybrid design ensures that S2 is invoked either for systematic exploration or as a robustness safeguard. The disjunctive combination yields an adaptive balance that outperforms either condition alone, as validated in our ablation study (Table 2). The meta-controller’s decisions form a closed loop system: high quality demonstrations from S2 are distilled into S1 via knowledge distillation (Appendix B.5.3), creating a virtuous cycle of policy improvement while progressively reducing long term reliance on costly deliberation.